OpenStudy (anonymous):

What is the sum of a 12–term arithmetic sequence where the last term is 13 and the common difference is –10?

OpenStudy (asnaseer):

ok, so here you are given \(n=12\), \(a_{12}=13\) and \(d=-10\)

OpenStudy (asnaseer):

now, do you recall the formula for calculating the n'th term of a sequence?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

you can use that formula to work out \(a_1\)

OpenStudy (asnaseer):

once you have that, then you can work out the sum \(S_{12}\)

OpenStudy (anonymous):

Is the first term 3?

OpenStudy (asnaseer):

no

OpenStudy (anonymous):

What formula do I use to find the first term?

OpenStudy (asnaseer):

ok, think about how would solve this problem. you are asked to find the sum of a 12 term series, and you of two formulas for the sum:\[S_n=\frac{n}{2}(a_1+a_n)\tag{1}\]\[S_n=\frac{n}{2}(2a_1+(n-1)d)\tag{2}\]you are given the values of \(n\), \(a_n\) and \(d\). so which of these two formulas would you pick to get the sum?

OpenStudy (asnaseer):

can you see that in both formulas you need to know the value of \(a_1\)?

OpenStudy (anonymous):

yes, so how would you find the first term?

OpenStudy (asnaseer):

good - so you are thinking now :) recall the formula for the n'th term:\[a_n=a_1+(n-1)d\tag{3}\]in this formula you know everything except \(a_1\) - which is exactly the value you need in order to use either formula (1) or (2) above. so you can first re-arrange formula (3) to get an expression for \(a_1\) as follows:\[a_1=a_n-(n-1)d\tag{4}\]so now you can use this formula to work out \(a_1\), and then use either formula (1) or (2) to work out the sum.

OpenStudy (anonymous):

133?

OpenStudy (asnaseer):

what is 133?

OpenStudy (anonymous):

first term?

OpenStudy (asnaseer):

can you please show your working?

OpenStudy (anonymous):

13 - 12(-10)

OpenStudy (asnaseer):

look carefully at the formula you are using

OpenStudy (asnaseer):

\(d\) is multiplied by \((n-1)\) not by \(n\)

OpenStudy (anonymous):

so it would be 11?

OpenStudy (anonymous):

13 - (11)(-10)

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

123?

OpenStudy (asnaseer):

yes - that is the first term

OpenStudy (anonymous):

final answer = 816?

OpenStudy (asnaseer):

perfect! well done again! :)